solve and check
. 8x – 11 = 2x + 1

Answer by 303795(602)

8x – 11 = 2x + 1
Add 11 to each side of the equation.
8x – 11 + 11 = 2x + 1 + 11
8x = 2x + 12
Subtract 2x from each side of the equation
8x - 2x = 2x + 12 - 2x
6x = 12
divide each side by 6 to get
x = 2
Substitute x = 2 into the original equation
8*2 – 11 = 2*2 + 1
16 - 11 = 4 + 1
5 = 5
ie the value x=2 is correct